Reports 1-1 of 1 Clear search Modify search
AdV-DET (Commissioning)
gouaty, magazzu - 11:21 Thursday 08 April 2021 (51352) Print this report
Recentering of B5 and B1p quadrants
This morning, Andrea set up the single bounce configuration (BS+NI aligned) and then we worked on the centering of the B5 and B1p quadrants on SDB2.

We took this opportunity to adjust the offsets in the H,V,Sum signals of the B5 quadrants.
Then we recentered the beam on the quadrants. The following adjustments were performed with picomotors:
B5_M1_H: -100 steps ; B5_M1_V: +175 steps ; B5_M2_H: -430 steps ; B5_M2_V: +430 steps
The galvo loops were then closed with corrections below 0.4V, as shown in Fig.1. The step visible in the Sum channels around 07h44 is actually due to the switch off of the picomotor drivers.
The same work was then performed on the B1p quadrants. The loops were initially closed with corrections around 1.5V. We reduced the corrections below 0.3V with the following picomotors adjustment:
B1p_M1_H: -225 steps ; B5_M1_V: +250 steps ; B5_M2_H: -200 steps ; B5_M2_V: -260 steps
We also performed a test of the 2f centering on the B1p quadrants between 07h58 and 08h01 (Fig.2). It worked but was much more noisier, probably because of the single bounce configuration.

After this realignment, we opened the galvo loops and close the shutters of the quadrants to keep the quadrants safe during the CITF realignment.
Images attached to this report
Search Help
×

Warning

×